# Likelihood-based Mitigation of Evaluation Bias in Large Language Models

### v2

@@ 1 Introduction
In this paper, we introduce likelihood bias, where LLM-based evaluators overrate high-likelihood sentences and underrate low-likelihood ones compared to human scores. Figure 1 shows one example of likelihood bias. Here, a biased evaluator gives a lower score of 3/5 to a correct but low-likelihood output (top) while giving a higher score of 5/5 to a high-likelihood output (bottom), based on the criteria of relevance. Addressing this issue, we propose a method that a) quantifies and b) mitigates likelihood bias. We quantify the bias by using the correlation between the disparity in evaluation scores generated by LLMs and those provided by human evaluators, and the likelihood of a target text. Our bias reduction method identifies and utilizes highly biased instances as few-shot examples for in-context learning.
The extent of likelihood bias may vary with the evaluation criteria used. For instance, likelihood bias is anticipated to be more pronounced in criteria like relevance, which is less directly related to likelihood. Conversely, the bias is expected to be less significant in criteria like fluency, which is closely related to likelihood. To verify this characteristic of likelihood bias, we adopt two tasks: data-to-text and Grammatical Error Correction (GEC). We use these tasks because, unlike most existing data Freitag et al. (2021); Guan et al. (2021); Kamalloo et al. (2023), the evaluation data for these tasks include multiple criteria such as fluency and relevance.
Our experimental results show that both evaluators based on GPT-3.5 and Llama2-13B Touvron et al. (2023) indeed suffer from likelihood bias. Moreover, our bias reduction method mitigates likelihood bias, and improves evaluation performance in many cases.
@@ 2.1 Measuring Likelihood Bias
When measuring the bias, we choose few-shot example instances at random.
@@ 2.2 Mitigating Likelihood Bias
Figure 2 plots LS against US in order to show the likelihood bias of multiple hypothetical evaluators. Each point represents a pair of scores for an instance. The BiasScore corresponds to the slope of the main cluster of instances.
@@ 3.2 Measuring Likelihood Bias
relevance .43 .28 .31 * .15 † .35 .25 .38 .23
fluency .26 .20 .29 .00 * .41 .33 .55 * .52 †
@@
data coverage .40 .24 .32 * .16 .20 .24 .30 † .25
total (micro) .38 .17 .32 † .02 † .48 .40 .58 * .46
GEC grammar .46 .24 .37 † .24 .48 .45 .54 .46
@@ 4 Conclusion
While we do not foresee any ethical risks caused by our research, LLMs not only exhibit biased likelihood based on surface-level information such as words and sentence structure but also on information like gender, religion, and race Kaneko et al. (2023); Oba et al. (2023); Anantaprayoon et al. (2023). For instance, LLMs might assign a higher likelihood to “She is a nurse” compared to “He is a nurse”. Reducing likelihood bias could potentially address social bias in evaluators. However, it is worth noting that this study does not investigate such aspects, and this remains a task for future research.
@@ Acknowledgements
These research results were obtained from the commissioned research (No.22501) by National Institute of Information and Communications Technology (NICT) , Japan. Also, this work was partially supported by JSPS KAKENHI Grant Number 19H01118. Finally, We are grateful to Simone Teufel, a professor at the University of Cambridge, for her valuable advice and feedback.

### v3

@@ 1 Introduction
In this paper, we introduce likelihood bias, where LLM-based evaluators overrate high-likelihood sentences and underrate low-likelihood ones compared to human scores. Figure 1 shows one example of likelihood bias. Here, a biased evaluator gives a lower score of 3/5 to a correct but low-likelihood output (top) while giving a higher score of 5/5 to a high-likelihood output (bottom), based on the criteria of data coverage. Addressing this issue, we propose a method that a) quantifies and b) mitigates likelihood bias. We quantify the bias by using the correlation between the disparity in evaluation scores generated by LLMs and those provided by human evaluators, and the likelihood of a target text. Our bias reduction method identifies and utilizes highly biased instances as few-shot examples for in-context learning.
The extent of likelihood bias may vary with the evaluation criteria used. For instance, likelihood bias is anticipated to be more pronounced in criteria like data coverage, which is less directly related to likelihood. Conversely, the bias is expected to be less significant in criteria like fluency, which is closely related to likelihood. To verify this characteristic of likelihood bias, we adopt two tasks: data-to-text and Grammatical Error Correction (GEC). We use these tasks because, unlike most existing data Freitag et al. (2021); Guan et al. (2021); Kamalloo et al. (2023), the evaluation data for these tasks include multiple criteria such as fluency, grammar, and data coverage.
Our experimental results show that both evaluators based on GPT-3.5 and Llama2-13B Touvron et al. (2023) indeed suffer from likelihood bias. Moreover, our bias reduction method mitigates likelihood bias, and improves evaluation performance in many cases. For reproducibility, we release our code  [^1] https://github.com/stjohn2007/likelihood_bias.
@@ 2.1 Measuring Likelihood Bias
When measuring the bias, we choose eight few-shot example instances at random.
@@ 2.2 Mitigating Likelihood Bias
Figure 2 plots LS against US in order to show the likelihood bias of multiple hypothetical evaluators  [^3] Please note that the figure is the pseudo-scatter plot that represents hypothetical evaluators. We do not include concrete values for LS and US in the figure since LS does not have the lower bound, and the scale of US depends on the dataset.. Each point represents a pair of scores for an instance. The BiasScore corresponds to the slope of the main cluster of instances.
@@ 3.2 Measuring Likelihood Bias
relevance .43 .28 .31 * ${}^{{\dagger}}$35 .25 .38 .23
fluency .26 .20 .29 .00 * .41 .33 .55 * ${}^{{\dagger}}$
@@
data coverage .40 .24 .32 * .16 .20 .24 ${}^{{\dagger}}$25
total (micro) .38 .17 ${}^{{\dagger}}$${}^{{\dagger}}$48 .40 .58 * .46
GEC grammar .46 .24 ${}^{{\dagger}}$24 .48 .45 .54 .46
@@ 3.3 Mitigating Likelihood Bias
4 Related Work
LLM-based evaluator
Kocmi and Federmann (2023) employ nine GPT models to evaluate translation tasks and verify that GPT-3.5 and larger models possess sufficient capabilities for task evaluation. Chiang and Lee (2023) show that several LLMs’ evaluation results are consistent with those obtained by expert human evaluation in open-ended story generation and adversarial attacks. They also find that LLM-based evaluators are stable over different prompts and sampling algorithms. Zheng et al. (2023) employ GPT-4 to evaluate the conversation ability of LLMs on two benchmarks: MT-bench and Chatbot Arena. Liu et al. (2023) propose G-EVAL, a framework of LLM-based evaluation. They incorporate chain-of-thoughts (CoT) Wei et al. (2023) and introduce a form-filling paradigm and scoring method for stable and fine-grained evaluation. Although LLM-based evaluators have been employed to evaluate several tasks, such as translation Kocmi et al. (2021), summarization Liu et al. (2023), story generation Chiang and Lee (2023), multi-turn conversation Zheng et al. (2023), in this work, we focus on two tasks, data-to-text and GEC, to inspect likelihood bias on multiple criteria.
Biases in LLM-based evaluators
Zheng et al. (2023) define self-enhancement bias as the tendency of LLM-based evaluators to favor the answers generated by themselves. Their preliminary experiments indicate the existence of this bias. A similar tendency is also reported by Liu et al. (2023). Also, Zheng et al. (2023) introduce verbosity bias, referring to the tendency of LLM-based evaluators to favor longer, more verbose responses. Saito et al. (2023) propose a metric to measure verbosity bias and find both GPT-4 and GPT-3.5 exhibit this bias according to their metric. In contrast to these findings, we introduce likelihood bias, where LLM-based evaluators evaluate high-likelihood sentences and underrate low-likelihood ones compared to human scores. Notably, this is the first work simultaneously proposing methodologies to quantify and mitigate a bias in LLM-based evaluators.
@@ 5 Conclusion
In future work, we aim to investigate the relationship between other biases in LLM-based evaluators, such as self-enhancement bias Zheng et al. (2023) and verbosity bias Zheng et al. (2023); Saito et al. (2023), and likelihood bias. For instance, we may be able to explain self-enhancement bias through the concept of likelihood bias, as the answers generated by LLMs might have high likelihoods calculated by the models themselves. Additionally, we plan to examine the impact of instruction-tuning and model size on likelihood bias.
@@
While we do not foresee any ethical risks caused by our research, LLMs not only exhibit biased likelihood based on surface-level information such as words and sentence structure but also on information like gender, religion, and race Kaneko et al. (2023); Oba et al. (2023); Anantaprayoon et al. (2023). For instance, LLMs might assign a higher likelihood to “She is a nurse” compared to “He is a nurse”. Reducing likelihood bias could potentially address social bias in evaluators. However, it is worth noting that this work does not investigate such aspects, and this remains a task for future research.
@@ Acknowledgements
These research results were obtained from the commissioned research (No.22501) by National Institute of Information and Communications Technology (NICT) , Japan. We are grateful to Simone Teufel, a professor at the University of Cambridge, for her valuable advice and feedback.